OpenStudy (anonymous):

Rewrite the square-root expression as an imaginary number. sqrt(-49)

OpenStudy (bradely):

sqrt(-49) =sqrt(49(-1)) -i=i^2 =sqrt(49(i^2)) =sqrt(49)*sqrt(i^2) =7i Source: http://www.mathskey.com/question2answer/

OpenStudy (anonymous):

Thanks!